Three charged marties are glued to a nonconducting surface and are placed in the dagram as shown. The charges of each marble are , -6.50 C, q 1.71 C, and a,- -2.37 C. Marble a, isa dstance,ー3.00 cm to the left of the martie av while martie 4, is adtance r -2.00 a to the right of the martie 42 asshown. Calculate the magnitude of the electric field a dsta cer-1.00 am to the lelt of the center marble N/C Another marble is placed 1 om to the lefht of the midde marble I this new marble has a charge of 3 magnitude direction .83 pc, calculate magnitude and direction of the force on it
Answer #1

E = Σ kQ / d²

At the point in question, the field due to q1 is positive (to the right), due to q2 is negative (to the left) and due to q3 is positive (to the right). Then

E = 8.99e9N·m²/C² * (6.50/(0.02m)² - 1.71/(0.01m)² + 2.37/(0.03m)²) * 1e-6C

E = 1.605e7 N/C ◄ field at 1 cm left of center marble

For fourth marble:

F = q*E = 3.68e-6C * 1.605e7N/C = 59.1 N ◄ [Verify the value of q as it is not clear in the image]

and the direction is "to the right." ◄